Given the following information for Mako Company please answer the following questions:

Sales

$500,000
Turnover

3

Margin

0.15

Desired Rate of Return

20%

What is the Residual Income?

Multiple Choice

  • $25,000.

  • $41,666

  • $66,667

  • $91,667
Answer #1

Margin=Net operating income/Sales

Net operating income=(500,000*15%)=$75000

Turnover=Sales/Average operating assets

Average operating assets=(500,000/3)=$166,666.67(Approx)

Residual income=Net operating income-(Average operating assets*Desired rate of return)

=75000-(166,666.67*20%)

=$41,666(Approx).